# XIFBench: Evaluating Large Language Models on Multilingual Instruction Following

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable instruction-following capabilities across various applications. However, their performance in multilingual settings remains poorly understood, as existing evaluations lack fine-grained constraint analysis. We introduce XIFBench, a comprehensive constraint-based benchmark for assessing multilingual instruction-following abilities of LLMs, featuring a novel taxonomy of five constraint categories and 465 parallel instructions across six languages spanning different resource levels. To ensure consistent cross-lingual evaluation, we develop a requirement-based protocol that leverages English requirements as semantic anchors. These requirements are then used to validate the translations across languages. Extensive experiments with various LLMs reveal notable variations in instruction-following performance across resource levels, identifying key influencing factors such as constraint categories, instruction complexity, and cultural specificity.  [^1] Our code and data will be released upon acceptance.
XIFBench: Evaluating Large Language Models on
Multilingual Instruction Following
Zhenyu Li1, Kehai Chen1,thanks:   Corresponding author., Yunfei Long2, Xuefeng Bai1, Yaoyin Zhang3, Xuchen Wei3, Juntao Li4, Min Zhang1 1Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China 2School of Computer Science and Electronic Engineering, University of Essex, Colchester, United Kingdom 3School of Computer Science and Technology, Harbin Institute of Technology, Shenzhen, China 4School of Computer Science and Technology, Soochow University, China 23s051032@stu.hit.edu.cn, chenkehai@hit.edu.cn, yl20051@essex.ac.uk
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in various natural language processing (NLP) tasks and real-world applications (Zhao et al., 2023). A key factor behind their widespread adoption is their ability to follow instructions (Ouyang et al., 2022; Wei et al., 2022), allowing them to comprehend and execute user commands while adhering to specified requirements (Zhou et al., 2023b; He et al., 2024b). Consequently, ensuring consistent multilingual instruction-following capability has attracted significant attention in advancing LLMs (Qin et al., 2024a; Shaham et al., 2024).
@@
Figure 1: Discrepancies in LLMs’ instruction following across languages. Given English instruction and Chinese translation, Llama-3.1-8B responses show varying capabilities in the following constraints, as seen in constraint-based evaluation. A concise back-translation (b.t.) of responses is provided for reference.
Despite the promise of multilingual instruction following, significant challenges remain due to disparities in language resources (Figure 1). Recent studies (Shaham et al., 2024; Weber et al., 2024) reveal substantial discrepancies in LLMs’ instruction-following capabilities across high- and low-resource languages. Most existing evaluations rely on simple, direct instructions and holistic assessment methods, such as pairwise comparison (Li et al., 2023) or direct scoring (Zheng et al., 2023), which fail to capture the complexity of real-world instructions and provide limited insights into performance gaps. Constraint-based evaluations have emerged as a promising alternative (Jiang et al., 2024; Qin et al., 2024b), as constraints encapsulate essential requirements that LLMs should satisfy. Constraints not only serve as fundamental components for instruction construction and evaluation but also enable a more fine-grained analysis of instruction-following capabilities in high-resource languages like English and Chinese. However, this leaves a critical gap in understanding how LLMs handle constraints across diverse languages, particularly for medium- and low-resource ones.
Motivated by this challenge, this paper aims to quantify and analyze multilingual instruction-following variations and their causes systematically. To achieve this, we introduce XIFBench, a comprehensive benchmark featuring three key innovations: (1) It is built upon a novel constraint taxonomy covering five categories—Content, Style, Situation, Format, and Numerical—with 21 dimensions carefully selected to ensure intra-category consistency and cross-lingual applicability by excluding language-dependent constraints. This taxonomy enables the automatic construction of instruction sets with balanced constraint categories and difficulty distributions. (2) To ensure consistent evaluation across languages, we develop a requirement-based protocol that decomposes complex instructions into evaluation requirements—English YES/NO questions serving as semantic anchors for cross-lingual assessment. This approach enables fine-grained, interpretable evaluation of instruction-following capabilities. (3) XIFBench provides high-quality parallel instruction sets in six languages, spanning high-resource (English, Chinese), medium-resource (Russian, Arabic), and low-resource (Hindi, Swahili) languages, with validation of requirement equivalence across translations. Extensive experiments with various LLMs reveal notable variations in instruction-following performance across resource levels, identifying key influencing factors such as constraint categories, instruction complexity, and cultural specificity.
@@ 2 Related Work
Instruction Following Evaluation
@@
As the fundamental capability of LLMs to align with human intents, InstructGPT (Ouyang et al., 2022) evaluated instruction following through human evaluation, employing direct scoring and binary adherence checks. Later works like AlpacaEval (Li et al., 2023) and MT-Bench (Zheng et al., 2023) utilized LLM judges for pairwise comparison and direct scoring, respectively. To assess LLMs on more complex tasks, WizardLM (Xu et al., 2024) and LIMA (Zhou et al., 2023a) introduced diverse and challenging instructions, although they still relied on holistic evaluation methods.
@@
The advancement of constraint-based evaluation has enabled fine-grained evaluation of instruction following. CELLO (He et al., 2024a) and CoDI-Eval (Chen et al., 2024) advanced this direction by collecting complex instructions and applying rule-based constraint checks. While IFEval (Zhou et al., 2023b) focused solely on rule-verifiable format constraints, FollowBench (Jiang et al., 2024) expanded to multiple constraint types (e.g., content, situation, style) and analyzed the impact of constraint quantity with LLM-based evaluation. InfoBench (Qin et al., 2024b) introduced requirement checklists for precise assessment and has been widely adopted in later studies. CFBench (Zhang et al., 2024a) developed a comprehensive constraint taxonomy from real-world scenarios, and ComplexBench (Wen et al., 2024) further investigated the impact of constraint composition types. However, most studies primarily focus on English and Chinese, leaving a gap in understanding instruction across languages.
Multilingual Instruction Following Evaluation
@@
Evaluating instruction following in multilingual contexts often relies on translated English-centric benchmarks like AlpacaEval, as demonstrated in various instruction tuning and evaluation studies (Zhang et al., 2024b; Li et al., 2024a; Shaham et al., 2024; Weber et al., 2024). Some works introduce language-specific benchmarks, such as Arabic MT-Bench (Boughorbel and Hawasly, 2023) and CIFBench (Li et al., 2024b) for Chinese. Although Aya Evaluation Suite (Singh et al., 2024) provides native annotations and translations across 101 languages, its evaluation approaches remain holistic.
Constraint-based evaluation remains underexplored in multilingual settings. A notable effort in this direction is Multi-IF (He et al., 2024b), which extends IFEval to multi-turn scenarios across multiple languages. However, Multi-IF primarily focuses on format constraints, leaving broader aspects of constraint adherence underexplored. In this paper, we propose XIFBench, an automatically constructed multilingual benchmark with a comprehensive constraint taxonomy for evaluating instruction following across resource levels.
Figure 2: The automated pipeline for constructing XIFBench, consisting of three stages with six steps: Constraint Augmentation (§3.1.2), Requirement Structuring (§3.1.3), and Multilingual Expansion (§3.1.4).
@@ 3.1.1 Instruction Preparation
Seed Instruction Selection
To ensure the diversity and representativeness of XIFBench, we source evaluation instructions from AlpacaEval (Li et al., 2023), WizardLM (Xu et al., 2024), and LIMA (Zhou et al., 2023a), which are widely used benchmarks for assessing multilingual instruction following capabilities (Zhang et al., 2024b; Li et al., 2024a; Shaham et al., 2024; Weber et al., 2024). To prevent data leakage, we include only evaluation sets and omit all training data. To promote task diversity, we perform hierarchical clustering on these instructions, resulting in 131 distinct clusters. We then select one representative instruction from each cluster, ensuring a diverse and comprehensive instruction set. See Appendix A.1.1 for details.
Instruction Filtering and Annotation
While existing benchmarks provide a useful starting point, not all instructions are suitable for multilingual evaluation or further augmentation. Thus, we manually filter out ambiguous, overly difficult, or language-dependent instructions (e.g., capitalized response, alliteration), yielding a refined set of 106 instructions, named the Easy Set. Each instruction is further annotated with cultural accessibility labels to indicate culturally specific references (e.g., Jesus Christ, YouTube) that may not be universally understood across languages. Additionally, we decompose each instruction into core instruction and input materials to facilitate subsequent automated construction. See Appendix A.1.2 for details.
@@ 3.1.2 Constraint Augmentation
Constraint Taxonomy
To systematically evaluate the impact of different constraint categories on multilingual instruction following, we propose a novel taxonomy comprising 5 categories and 21 dimensions, inspired by prior constraint-based evaluations (Jiang et al., 2024; Qin et al., 2024b; Zhang et al., 2024a). As illustrated in Figure 2, each constraint follows the structure category - dimension - specification, such as content - include content - discuss scientific research, guiding LLMs to include scientific research in their responses.
Our taxonomy covers five major categories: (1) Content, which specifies what information should be included or excluded; (2) Style, which defines the tone and writing manners; (3) Situation, which describes contextual settings such as assumed roles or environments; (4) Format, which outlines structural requirements for responses; and (5) Numerical, which involves quantitative factors like length or item counts. To ensure cross-lingual applicability, we intentionally exclude Linguistic constraints that depend on language-specific properties (e.g., uppercase, alliteration, word count), which were a focus in prior English- and Chinese-centric taxonomies. See Appendix A.2.1 for details.
Constraint Brainstorming
For each core instruction, we prompt GPT-4o (OpenAI, 2024) [^2] We use gpt-4o-2024-08-06 for all constructions. to generate multiple independent constraints across all 5 categories and 21 dimensions, ensuring that each dimension is adequately covered. This process results in a balanced constraint set while maximizing the variety of applicable constraints for each instruction. Additionally, it allows precise control over the types of constraints introduced in subsequent steps. See Appendix A.2.2 for details.
Constraint Combination
To balance instruction difficulty and diversity, we sample constraints from the brainstormed set to construct instructions with 1 to 5 additional constraints. The sampling algorithm prioritizes underrepresented categories for even distribution. We then prompt GPT-4o to integrate the selected constraints into fluent and natural instruction. This process yields the Hard Set, containing 530 combined instructions. See Appendix A.2.3 for details.
Instruction Validation
To ensure instruction quality, two human annotators independently assess each instruction for clarity, cultural accessibility, and linguistic dependency (i.e., whether it involves linguistic constraints). Instructions with defects are removed along with all their variations in both the Easy Set and Hard Set to maintain consistency. This validation yields 93 high-quality instructions in the Easy Set and 465 in the Hard Set. See Appendix A.2.4 for details.
@@ 3.1.3 Requirement Structuring
Requirement Decomposition
To enable fine-grained evaluation, we decompose instructions into atomic requirements, following prior works Qin et al. (2024b); Zhang et al. (2024a); Wen et al. (2024). Moreover, we automatically complete this process by prompting GPT-4o to extract evaluation requirements from the instructions in both Easy Set and Hard Set. As shown in Figure 2, each requirement is a binary (YES/NO) question (e.g., Is each section limited to 2 sentences?), ensuring a precise and interpretable assessment. See Appendix A.3.1 for details.
Requirement Categorization
Since the initial core instructions contain inherent constraints, and constraints from the combination step may be rephrased differently, we need to map the extracted evaluation requirements back to their original categories. To achieve this, we prompt GPT-4o to classify each requirement into the most suitable category and dimension within our predefined taxonomy. This automated process ensures consistency and facilitates a detailed analysis of the adherence to constraints in languages and models. See Appendix A.3.2 for details.
Requirement Assessment
To verify the accuracy of requirement extraction and categorization, we sample 10% of instructions (10 from the Easy Set, 50 from the Hard Set) and ask two annotators to independently evaluate the extracted requirements on four dimensions: (1) Explicitness: The requirement must be explicitly stated in the instruction. (2) Completeness: It must cover all explicitly stated constraints. (3) Atomicity: It should be atomic, addressing a single aspect. (4) Categorization: It should be assigned to the correct category. Our evaluation confirms its high reliability. See Appendix A.3.3 for details.
@@ 3.1.4 Multilingual Expansion
Language Selection
To ensure broad multilingual coverage while maintaining feasibility, we select six representative languages—English, Chinese, Russian, Arabic, Hindi, and Swahili—based on prior multilingual instruction tuning works (Shaham et al., 2024; Li et al., 2024a). These languages span different resource levels: high (English, Chinese), medium (Russian, Arabic), and low (Hindi, Swahili), following Joshi et al. (2020). This selection enables a systematic analysis of language resource impact on instruction following.
Parallel Translation
We translate the core instruction of each English instruction in Easy Set and the combined instruction in Hard Set into five other languages using Google Translate, while keeping the input materials unchanged. This ensures that both code snippets and structured formats remain consistent across languages.
Translation Validation
While automatic translation is generally reliable, subtle semantic shifts may still affect constraint-based evaluation. As shown in Figure 2, we use GPT-4o and Google Translate’s back-translation to verify requirement equivalence across languages, ensuring each translated instruction segment aligns with the original English version. Our analysis reveals that only 0.29% of evaluation requirements exhibit inconsistencies across languages, demonstrating sufficient translation quality for multilingual evaluation. See Appendix A.4 for details.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable instruction-following capabilities across various applications. However, their performance in multilingual settings lacks systematic investigation, with existing evaluations lacking fine-grained constraint analysis across diverse linguistic contexts. We introduce XIFBench, a comprehensive constraint-based benchmark for evaluating multilingual instruction-following abilities of LLMs, comprising 558 instructions with 0-5 additional constraints across five categories (Content, Style, Situation, Format, and Numerical) in six languages spanning different resource levels. To support reliable and consistent cross-lingual evaluation, we implement three methodological innovations: cultural accessibility annotation, constraint-level translation validation, and requirement-based evaluation using English requirements as semantic anchors across languages. Extensive experiments with various LLMs not only quantify performance disparities across resource levels but also provide detailed insights into how language resources, constraint categories, instruction complexity, and cultural specificity influence multilingual instruction-following. Our code and data are available at https://github.com/zhenyuli801/XIFBench.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in various natural language processing (NLP) tasks and real-world applications [1]. A significant driver of their widespread adoption is their instruction-following capability [2, 3], enabling them to understand and execute user commands while adhering to specified requirements [4, 5]. To support global applications and serve diverse linguistic communities, ensuring consistent multilingual instruction-following capability has attracted significant attention in advancing LLMs [6, 7].
@@
Figure 1: Discrepancies in LLMs’ instruction-following across languages. Given English instruction and Chinese translation, Llama-3.1-8B response exhibits varied constraint-following, as shown in constraint-based evaluation. A concise back-translation (b.t.) of response is provided for reference.
While LLMs have shown promise in multilingual instruction-following, recent studies [7, 8] reveal significant performance disparities across languages (Figure 1). Factors such as the composition of multilingual instruction data have been investigated as potential contributors to this variation. However, existing evaluations often involve relatively simple, direct instructions and coarse-grained evaluation methods, such as pairwise comparison [9] or direct scoring [10]. These approaches fall short in capturing the complexity of real-world instructions and offer limited insight into how instruction-inherent factors affect cross-lingual performance.
Constraint-based evaluations have emerged as a promising alternative for assessing instruction-following [11, 12], as constraints are essential requirements within user instructions that LLMs should satisfy. Constraints provides explicit targets for instruction formulation and evaluation, enabling a more fine-grained analysis of instruction-following capabilities. However, existing constraint-based frameworks are largely limited to high-resource languages like English and Chinese. This leaves a critical gap in understanding how multilingual properties, such as resource availability, influence LLMs’ constraint-following ability in diverse linguistic contexts.
To bridge these critical gaps in multilingual instruction-following evaluation, we introduce XIFBench, a comprehensive benchmark designed to assess LLMs’ capabilities with constraint-rich instructions across diverse linguistic contexts. XIFBench comprises 558 instructions, systematically augmented with 0-5 constraints drawn from a taxonomy covering five categories (Content, Style, Situation, Format, and Numerical) and 21 fine-grained dimensions. To ensure cross-lingual applicability, we specifically exclude language-dependent constraints. Beyond English, we translate these instructions into five languages (Chinese, Russian, Arabic, Hindi, and Swahili) representing high-, medium-, and low-resource settings. To enable insightful and reliable multilingual evaluation, we propose three methodological innovations: (1) We annotate the cultural accessibility of each instruction to investigate how these nuanced semantic elements impact cross-lingual instruction adherence, a factor often overlooked in existing evaluations. (2) We implement a constraint-level translation validation process that evaluates the semantic preservation of each constraint across languages. This ensures high fidelity and strict parallelism of constraints across languages. (3) We adapt the requirement-based evaluation methodology to multilingual contexts, using shared English requirements as semantic anchors. This approach ensures reliable and consistent instruction-following assessment across languages. Extensive experiments conducted on XIFBench with multiple LLMs not only quantify performance disparities across resource levels but also provide detailed, fine-grained insights into how factors such as language resources, constraint categories, instruction complexity, and cultural specificity influence multilingual instruction-following.
@@ 2 Related Work
Instruction-Following Evaluation
@@
As the fundamental capability of LLMs to align with human intents, InstructGPT [2] evaluated instruction-following through human evaluation, employing direct scoring and binary adherence checks. Later works like AlpacaEval [9] and MT-Bench [10] utilized LLM judges for pairwise comparison and direct scoring, respectively. To assess LLMs on more complex tasks, WizardLM [13] and LIMA [14] introduced diverse and challenging instructions, but they still relied on coarse-grained evaluation methods.
@@
The advancement of constraint-based evaluation has enabled fine-grained evaluation of instruction-following. CELLO [15] and CoDI-Eval [16] advanced this direction by collecting complex instructions and applying rule-based constraint checks. While IFEval [4] centers on objectively verifiable constraints, it predominantly covers format, numerical, and linguistic constraints. In contrast, FollowBench [11] expanded to more constraint types (e.g., content, situation, style) and analyzed the impact of constraint quantity with LLM-based evaluation. InfoBench [12] introduced requirement checklists for precise assessment and has been widely adopted in later studies. CFBench [rw:cf-bench] developed a comprehensive constraint taxonomy from real-world scenarios, and ComplexBench [17] further investigated the impact of constraint composition types. However, most studies primarily focus on English and Chinese, leaving a gap in understanding instruction-following across languages.
Multilingual Instruction-Following Evaluation
@@
To evaluate multilingual instruction-following capabilities, existing work has primarily relied on translated versions of English-centric benchmarks, such as AlpacaEval [9], as demonstrated in various instruction tuning and evaluation efforts [18, 19, 7, 8]. Some research introduce language-specific benchmarks, such as Arabic MT-Bench [20] and CIFBench [21] for Chinese. Although Aya Evaluation Suite [22] provides native annotations and translations across 101 languages, its evaluation approaches remain coarse-grained.
While constraint-based evaluation is well-established for English, its multilingual exploration remains limited. Recent efforts have extended IFEval’s framework to additional languages, like M-IFEval [23] covering English, French, Japanese and Spanish, and Multi-IF [5] extending to eight primarily Indo-European languages in multi-turn scenarios. However, these efforts largely target high- and medium-resource languages, thus offering limited insight into instruction-following capabilities across the full resource spectrum. Furthermore, by inheriting IFEval’s focus on objective constraints (e.g., format, numerical), these benchmarks may fail to evaluate multilingual instruction-following in scenarios involving semantically rich constraints (e.g., style, situation), which are common in real-world instructions and potentially more sensitive to cross-lingual variations. In contrast, our work examines multilingual instruction-following across 6 languages spanning high, medium, and low resources, while including a comprehensive taxonomy of both objective and semantic constraints.
@@ 3.1.1 Instruction Preparation
Seed Instruction Selection  To ensure the diversity and representativeness of XIFBench, we source evaluation instructions from AlpacaEval [9], WizardLM [13], and LIMA [14], which are widely used benchmarks for assessing multilingual instruction following capabilities [18, 19, 7, 8]. To prevent data leakage, we include only evaluation sets and omit all training data. To promote task diversity, we perform hierarchical clustering on these instructions, resulting in 131 distinct clusters. We then select one representative instruction from each cluster, ensuring a diverse and comprehensive instruction set. See Appendix C.1.1 for details.
Instruction Filtering and Annotation  While existing benchmarks provide a useful starting point, not all instructions are suitable for multilingual evaluation or further augmentation. Thus, we manually filter out ambiguous, overly difficult, or language-dependent instructions (e.g., capitalized response, alliteration), yielding a refined set of 106 instructions, named the Easy Set. Each instruction is further annotated with cultural accessibility labels to indicate culturally specific references (e.g., Jesus Christ, YouTube) that may not be universally understood across languages. Additionally, we decompose each instruction into core instruction and input materials to facilitate subsequent automated construction. See Appendix C.1.2 for details.
Figure 2: The automated pipeline for constructing XIFBench, consisting of three stages with six steps: Constraint Augmentation (§3.1.2), Requirement Structuring (§3.1.3), and Multilingual Expansion (§3.1.4). The example shown follows the same instruction as in Figure 1.
@@ 3.1.2 Constraint Augmentation
Constraint Taxonomy  To systematically evaluate the impact of different constraint categories on multilingual instruction-following, we construct a taxonomy of 5 categories and 21 dimensions, building upon insights from prior constraint-based evaluations [11, 12, rw:cf-bench]. As illustrated in Figure 2, each constraint follows the structure category - dimension - specification, such as content - include content - discuss scientific research, guiding LLMs to include scientific research in their responses.
Our taxonomy covers five categories: (1) Content, which specifies what information should be included; (2) Style, which defines the tone and writing manners; (3) Situation, which describes contextual settings such as assumed roles or environments; (4) Format, which outlines structural requirements for responses; and (5) Numerical, which involves quantitative factors like length or item counts. To ensure cross-lingual applicability, we intentionally exclude language-dependent constraints that depend on language-specific properties (e.g., uppercase, alliteration, word count), which were a focus in prior English- and Chinese-centric taxonomies. See Appendix C.2.1 for details.
Constraint Brainstorming  For each core instruction, we prompt GPT-4o [24] [^1] We use gpt-4o-2024-08-06 for all constructions. to generate multiple independent constraints across all 5 categories and 21 dimensions, ensuring that each dimension is adequately covered. This process results in a balanced constraint set while maximizing the variety of applicable constraints for each instruction. Additionally, it allows precise control over the types of constraints introduced in subsequent steps. See Appendix C.2.2 for details.
Constraint Combination  To balance instruction difficulty and diversity, we sample constraints from the brainstormed set to construct instructions with 1 to 5 additional constraints. The sampling algorithm prioritizes underrepresented categories for even distribution. We then prompt GPT-4o to integrate the selected constraints into fluent and natural instruction. This process yields the Hard Set, containing 530 combined instructions. See Appendix C.2.3 for details.
Instruction Validation  To ensure instruction quality, two human annotators independently assess each instruction in the Hard Set for clarity and linguistic dependency (i.e., whether it involves language-dependent constraints), and annotate its cultural accessibility. Instructions with defects and all their variations are removed, ensuring each Easy Set instruction has exactly five Hard Set counterparts with 1-5 added constraints. This validation yields 93 high-quality instructions in the Easy Set and 465 in the Hard Set. See Appendix C.2.4 for details.
@@ 3.1.3 Requirement Structuring
Requirement Decomposition  To enable fine-grained evaluation, and because these constraints are compound or altered during integration, we decompose instructions into atomic requirements, following prior works [12, rw:cf-bench, 17]. We complete this process by prompting GPT-4o to extract evaluation requirements from instructions in both Easy Set and Hard Set. As shown in Figure 2, each requirement is a binary (YES/NO) question (e.g., Is each section limited to 2 sentences?), ensuring a precise and interpretable assessment. See Appendix C.3.1 for details.
Requirement Categorization  Since the initial core instructions contain inherent constraints, and constraints from the combination step may be rephrased differently, we need to map the extracted evaluation requirements back to their original categories. To achieve this, we prompt GPT-4o to classify each requirement into the most suitable category and dimension within our predefined taxonomy. This automated process ensures consistency and facilitates a detailed analysis of the adherence to constraints in languages and models. See Appendix C.3.2 for details.
Requirement Assessment  To verify the accuracy of requirement extraction and categorization, we sample 10% of instructions (10 from the Easy Set, 50 from the Hard Set) and ask two annotators to independently evaluate the extracted requirements on four dimensions: (1) Explicitness: The requirement must be explicitly stated in the instruction. (2) Completeness: It must cover all explicitly stated constraints. (3) Atomicity: It should be atomic, addressing a single aspect. (4) Categorization: It should be assigned to the correct category. Our evaluation confirms their high quality, with at least 93.3% of requirements adhering to each criterion. See Appendix C.3.3 for details.
@@ 3.1.4 Multilingual Expansion
Language Selection  To ensure broad multilingual coverage while balancing evaluation costs, we select six representative languages—English, Chinese, Russian, Arabic, Hindi, and Swahili—based on prior multilingual instruction tuning works [7, 19]. These languages span different resource levels: high (English, Chinese), medium (Russian, Arabic), and low (Hindi, Swahili), following Joshi et al. [25]. The selection encompasses Indo-European (English, Russian, Hindi), Sino-Tibetan (Chinese), Afro-Asiatic (Arabic), and Niger-Congo (Swahili) language families, written in Latin, Cyrillic, Chinese, Arabic, and Devanagari scripts. This selection enables a systematic analysis of language resource impact on instruction following across diverse linguistic contexts.
Parallel Translation  We translate the core instruction of each English instruction in Easy Set and the combined instruction in Hard Set into five other languages using Google Translate, while keeping the input materials unchanged. This ensures that both code snippets and structured formats remain consistent across languages.
Translation Validation  While automatic translation is generally reliable, subtle semantic shifts can affect constraint-based evaluation. For example, the Chinese translation of "use metaphors for bacterial movement" in Figure 1 more closely resembles "use bacterial movement’s metaphors." To validate such issues, as illustrated in Figure 2, we use GPT-4o and Google Translate’s back-translation to evaluate the semantic preservation of each evaluation requirement across languages, ensuring that the translated instruction segment for each requirement aligns with its English counterpart. The aforementioned discrepancy is successfully identified and rated as 3/5, indicating borderline preservation. Our analysis shows that less than 1.4% of evaluation requirements exhibit inconsistencies across languages. Human annotations on a sampled subset confirm this trend and show moderate agreement with automatic scores, confirming that the translation quality is sufficient for reliable multilingual evaluation. See Appendix C.4 for details.